OpenStudy (anonymous):

What is the definite integral of arctan(1/x) evaluated from 1 to sqrt(3)?

OpenStudy (anonymous):

Integration by parts seems to be the way to go as arctan will become nicer upon differentiation

OpenStudy (anonymous):

satellite73 please help me

OpenStudy (anonymous):

\[\begin{matrix}u=\arctan\left(\frac{1}{x}\right)&dv=dx\\ du=\frac{-\frac{1}{x^2}}{\frac{1}{x^2}+1}dx&v=x\end{matrix}\] \[x\arctan\left(\frac{1}{x}\right)-\int x\frac{-\frac{1}{x^2}}{\frac{1}{x^2}+1}dx\\ x\arctan\left(\frac{1}{x}\right)+\int \frac{\frac{1}{x}}{\frac{1}{x^2}+1}dx\] This integral seems tricky at first... Not sure if there's a closed form.

OpenStudy (anonymous):

\[x\arctan\left(\frac{1}{x}\right)-\int\frac{1}{\left(\frac{1}{x^2}+1\right)x}dx\] That might help a bit.

OpenStudy (anonymous):

okk thanks :)

zepdrix (zepdrix):

Multiplying that last integral \(\large \int \dfrac{\frac{1}{x}}{\frac{1}{x^2}+1}dx\) by \(\large \dfrac{x^2}{x^2}\) should clean it up a little bit. :)